OpenStudy (anonymous):

help me plz what is the edge of a cube that has 10cm^3

OpenStudy (anonymous):

Your question makes no sense.

OpenStudy (anonymous):

He wants to know the length of one edge... The volume of a cube is given by\[V = a^3 = a*a*a\]where a is the length of one edge.. You can do the rest yourself ;)

OpenStudy (anonymous):

I guess 10 cm^3 is the volume of the cube. And Volume=a*a*a, where a is the length of one edge of cube. As we know all sides are equal in a cube, Therefore Volume=a^3 Now Volume=10 cm^3 Hence, a^3=10 Therefore,\[a^3=10\] and \[a=\sqrt[3]10\] and that comes out to be 2.15443469003. Therefore the edge of the cube is approx. 2.15 cm.